Agenda/Minutes on 2020/12/25
CRY+VIS Interface Meeting
Participants: Takahashi, Sato, Aso, Miyo, Ushiba, Tomaru, Ikeda, Kimura
- Schedule for IX: Connection of ITMX was scheduled on 2/8. Cryostat will be closed after that.
- Schedule for IY: Cooling of the shield is on going. One week delay. Cooling of payload is scheduled from 1/12. Test of moving mass, test of accelerometer, and check of window are planed after that.
- Corrugate tube fixing jig: The height of top chamber went down by 4.4mm after evacuation. Still under the investigation.
- Flow chart: Two-step commissioning in vacuum was approved.
- Acceptance check items: Updated document for cryo-payload.
- Jig for FC painting: Ordered.
PF actuator: Measurement of coefficients --> calculation of the noise (by the end of Jan.).
Participants: Takahashi, Tanaka, Fabian, Sato, Aso, Miyo
- Preparation for IXV.
- Blade bending jig was moved to the IXV booth.
- LVDT coils were assembled onto the ACCs.
- Horiuchi-san (MPC) will join to the Type-A team from 1/5.
Week of the 14th of December
- Yano-san repaired a damaged thread in a BS LVDT component. He also compiled the bill is screws.
- Hirata-san and Yano-san went to wash the BS LVDT components ot the University of Toyama.
- The components were brought back to the tunnel on Friday.
- The plastic box was opened until Monday to let the components warm up to room temperature and thus avoid condensation.
Yano-san finished writing a nice procedure to assemble the LVDTs.
After receiving and washing new screws, we mounted the new lighter F0 keystone assembly top disc in SR3 (klog 15664). We removed 96 grams to compensate for the extra weight of the F0 yaw stepper motor mechanism.
- I finished writing the procedue for adjustting ballas mass in SR3. Hirata-san and Yano-san read it and they had comments.
- The temperature in BS area this week was still high (around 23.6 degrees). Yokozawa-san had misread my message and he thought such temperature was the one requested. Now he is aware of the temperature we want (close to 23.1 degrees).
Despite the temperature being high, we added ballast mass on top of BS BF in order to assess how easy or hard was going to be to keep the BF current balance (klog 15651 and comments):
- As planed, we put equal amounts of mass on geometrically opposite places: 66 grams in each place.
- It worked well in the sense that we didn't disturb the aligmment of the IMR and OSEMs. The largest change was -41 urad in IM-R, which is rather small for this type of adjustment.
- Transfer function measurements showed a healthy system.
- F1 fishing rod had been put in mid-range beforehand and the additional ballast moved keystone position down by 560 um to 423 um. We will need to add more ballast later because we aim to 212 um.
- The yaw measured by the oplev changed also, but this is due to the change in F1 height and coupling with cables.
Week of the 21st of December
- On Monday we receive the BS IP LVDT yokes:
- They were cleaned during 5 minuted in the ultrasonic bath with acetone in ordert to remove residual rust, likely coming out from threaded holes.
- We assembled completely two LVDT-actuator units; the other one was partially aseembled because we only have two jigs. After installing the first one we will finish the assemby of the third one.
- After assembly we checked the resistances of the coils to make sure there was no damage.
- We finished the assembly and installtion of the BS limit switches for the IP horizontal fishing rods.
- We attached extensions to some of the cables.
- Because of the lack of cable clamps in the designed we used inserted M8 vented SDC screws in nearby holes and used them as poles. We tied cables to them using kapton coated wire.
- We tested the switches manually and checked the medm screen gave the correct signal.
- In SRM we corrected the polarity of the F0 FR stepper motor.
- There was some troubleshooting to do: a cable broke at the attachment point to a pin (pin 2) and we fixed it.
- Using a spare stepper motor we checked the health of the unused channels in the SRM GAS Filter driver: two out of three work.
- Cable swapping for LVDT in PRM is necessary.
- IMMT2 modification:
Since there is little benefit to using IMMT2 transmitted light, the policy has been changed to dump them. Therefore, the change in IMMT2 frame was cancelled. japanese.txt
- Realtime model modification is on going.
[Takahashi] Tested the new LVDT combiner circuit (klog)
- [Tanaka] Designed modified circuit with wide bandpass filter for the secondary coil and gentle lowpass filter for actuator.
- Type-B remedy work is on going.
- Remedy work in IXV will start on 1/5.
- Procedure documents.
- Criteria in the acceptance test.
- SR3 suspension removal is not necessary. We will ask MPC about opening the hatch in SRM.
The week from 12/28
The week from 1/4
- Takahashi: 1/4-8
- Sato: 1/4-8
- Hirata: 1/4-8
- Yano: 1/4-8
- On 2020/1/8(Fri)